Hair Care Inc. is a wholesaler of hair supplies. Hair Care uses a perpetual inventory system. The following transactions (sum2. Compute the gross profit percentage. (Round your answer to the nearest whole percent.) Gross profit percentage %Journal entry worksheet 2 3 4 5 6 7 8 Record the sale of merchandise. Note: Enter debits before credits. Transaction GeneralJournal entry worksheet 3 4 5 6 7 8 Record the cost of merchandise sold. Note: Enter debits before credits. General Journal DJournal entry worksheet < 1 4 5 6 7 8 Record the merchandise received from the customers as unsatisfactory (but in perfect coJournal entry worksheet < 1 2 3 4 5 6 7 8 Record the original cost of merchandise returned. Note: Enter debits before creditsJournal entry worksheet < 1 2 3 4 6 7 8 Record the sale of merchandise. Note: Enter debits before credits. Transaction GeneraJournal entry worksheet < 1 2 3 4 5 6 7 Record the cost of merchandise sold. Note: Enter debits before credits. General JournJournal entry worksheet < 1 2 3 4 5 6 Record the collection of half of the balance owed by the customer in (c), paid within tJournal entry worksheet Record the allowance granted to the customer in (c). Note: Enter debits before credits. Transaction G4. Hair Care is considering a contract to sell merchandise to a hair salon chain for $14,700. This merchandise will cost Hair
Answer #1
Account Titles Debit Credit
a Cash $       50,900
      Sales Revenue $       50,900
Cost of Goods Sold $       28,497
      Inventory $       28,497
b Sales returns and allowances $             570
      Cash $             570
Inventory $             330
      Cost of Goods Sold $             330
c Accounts receivable $       10,000
      Sales Revenue $       10,000
Cost of Goods Sold $         4,600
      Inventory $         4,600
d Cash $         4,850
Sales Discount $             150
      Accounts Receivable $         4,850
e Sales returns and allowances $             145
      Accounts Receivable $             145
Sales Revenue $       60,900 =50900+10000
Net Sales $       60,035 =60900-570-150-145
Cost of Goods Sold $       32,767 =28497-330+4600
Gross Profit $       27,268
Gross Profit (%) 45.42%

Gross Profit (%) = Gross Profit / Net Sales x 100

Net Sales $       74,735 =60035+14700
Cost of Goods Sold $       42,767 =32767+10000
Gross Profit $       31,968
Gross Profit (%) 42.78%

It decreases Gross Profit Percentage